OpenStudy (anonymous):

question is 9y^2-55y+6/9y^2+180y-9 multiplyed by 6y^2+53y-9/54y^2-15y+1

OpenStudy (anonymous):

It's actuallly 9y^2+80-9 not 180

OpenStudy (anonymous):

\[= ((9y-1)(y-6)/(9y-1)(y+9)) x ((6y-1)(y+9)/(6y-1)(9y-1))\]\[= y-6\]

OpenStudy (anonymous):

sorry! =[(9y-1)(y-6)/(9y-1)(y+9)] x [(6y-1)(y+9)/(6y-1)(9y-1)] = y - 6/ 9y - 1 ^_^

OpenStudy (anonymous):

ugh!! Thank you soo much, I have been stuck on this one!!

OpenStudy (anonymous):

np :)